Sunlight strikes a piece of crown glass at an angl
Answer #1

We know that n is the index of refraction. and n of air= 1, so, n1 =1.

n1*sin(theta 1)= n2*sin(theta 2)

so for Red: 1*sin (30.2)=1.52* sin (theta 2)
so 0.503=1.52*sin (theta 2)
sin(theta)=0.3309

So, Theta = 19.32 degrees for Red


for blue same thing: 1*sin(30.2)=1.531*sin(theta 2)
0.503=1.531*sin (theta 2)
sin(theta 2)= 0.5793/1.531
sin(theta 2)=0.328

theta for blue = 19.18 degrees


Red - blue =19.32-19.18=0.14


Difference in angle of refraction is 0.14

b) For Red Ray,

sin( critical angle ) = n2/n1 = 1/1.520, therefore angle = 41.4 deg